Three sisters, Mimi, Alison and Tracy decide to buy a house. Mimi pays

1

_

3

of the cost, Alisa pays

2

_

5

of

the cost and Tracy pays the rest.


2.2.1 What fraction of the cost of the house does Tracy pay? [2]




1
Expert's answer
2020-07-02T04:44:32-0400

Mimi pays 1/3, Alisa pays 2/5

they both pay 1/3+2/5=(5+6)/15

Tracy pays the remaining fraction, 1-11/15=4/15

Tracy pays 4/15 of the cost of the house.